# A Generative Modeling / Physics-Informed Neural Network Approach to Random Differential Equations

### v1

@@ 1.1 Problem Formulation and Approach
$\begin{split}&\mathscr{A}_{\xi}u(\cdot,\xi)=\mathscr{F}_{\xi}\,,\end{split}$ (1)
where $\mathscr{A}_{\xi}$ is a $\xi-$parametrised differential operator (incorporating possible initial and boundary conditions), and $\mathscr{F}_{\xi}$ is the source. For each fixed $\xi$, $u(\cdot;\xi)$ denotes the solution of the differential equation, which we interpret as a mapping $U:\Xi\to\mathcal{H}$ with $U(\xi):=u(\cdot,\xi)$. The space $\mathcal{H}$ where $u$ takes values for each $\xi$ is an (infinite dimensional) Hilbert or Sobolev space. We assume that the random parameter $\xi\in\Xi$ is distributed according to a known Borel probability measure $\gamma$ on $\Xi$. Our aim is to approximate the random variable $U$ and its law $\nu_{u}\in\mathcal{P}(\mathcal{H})\,;$ here $\mathcal{P}(\mathscr{X})$ denotes the set of Borel probability measures on $\mathscr{X}\,.$
@@
One of the key successes of neural network methods is their ability to effectively approximate probability distributions. In fact, several important problems of interest rely on generating a probability distribution when the available data represent specific instances of the unknown distribution. Typically, if the data is represented by an empirical measure $\nu_{\,\text{Dat}}\in\mathcal{P}(\mathbb{R}^{k})$, we would like to generate an approximation of $\nu_{\,\text{True}}\in\mathcal{P}(\mathbb{R}^{k})$ by optimising
$\min_{\mu\in V_{\Theta}}d_{\mathscr{M}}\left(\mu,\nu_{\,\text{Dat}}\right)\,,$ (2)
@@
$\mu_{\theta^{*}}\cong\nu_{\,\text{True}},\qquad\text{where }\mu_{\theta^{*}}\in\operatorname*{arg\,min}_{\mu\in V_{\Theta}}d_{\mathscr{M}}\left(\mu,\nu_{\,\text{Dat}}\right)\,.$ (3)
Representative works utilising the aforementioned generative model framework for a variety of applications include [adler2018banach, goodfellow2014generative, arjovsky2017wassersteingan, zhang2018wassersteinganmodeltotal, creswell2018generative, jospin2022hands], and the references in [Katsoulakis_z2023meanfieldlab]. For a recent modelling perspective that interprets various generative models as mean field games, see [Katsoulakis_z2023meanfieldlab].
The above framework relies on various alternative choices, each one yielding fundamentally distinct algorithms at the end. We would like to explore the potential extension of this approach to approximate solutions of (1). To accomplish this we must first systematically construct discrete spaces $V_{\Theta}$ such that $V_{\Theta}\subset\mathcal{P}(\mathcal{H}).$ We adopt the following alternative approaches:
@@
$\mathscr{A}\odot\mu,\qquad\mu\in V_{\Theta}\subset\mathcal{P}(\mathcal{H})\,,$ (4)
@@
$\min_{\mu\in V_{\Theta}}d_{\tilde{\mathscr{M}}}\left(\mathscr{A}\odot\mu,\nu_{\,\mathscr{F}}\right)\,,$ (5)
where $\nu_{\,\mathscr{F}}\in\mathcal{P}(\mathcal{H}^{\text{imag}})$ is the distribution corresponding to the source $\mathscr{F}_{\xi},$ and $d_{\tilde{\mathscr{M}}}\left(\cdot,\cdot\right)$ a specified distance (or divergence) on $\mathcal{P}(\mathcal{H}^{\text{imag}})$.
@@
This work is structured to systematically develop and analyse a generative modelling approach for Uncertainty Quantification using Physics-Informed Neural Networks (PINNs). In Section 2, we introduce the underlying problem, and the corresponding probabilistic framework essential to our study. Next, we introduce new notational conventions and the variational formulation used. The variational principle forms the backbone of our approach, guiding the optimisation process for learning the solution distribution of the random differential equation. Section 3 presents the model construction in detail. We define a fundamental representation of Neural Measures tailored to the problem’s probabilistic nature and explore three distinct architectural frameworks, namely fully discrete randomised neural networks, Polynomial Chaos Expansion Neural Networks (PCE-NN), Galerkin-based neural networks. In Section 4, we shift focus to the variational distances used during training, with particular emphasis on the Wasserstein distance and its interpretation in terms of Dirac measures, which naturally relate it to $L^{p}$-type loss functionals. We also discuss alternative metrics and divergences that can be integrated into the framework. Section 5 presents numerical experiments that validate the proposed framework across three representative problems: a bistable ODE, a linear diffusion PDE, and a reaction-diffusion PDE. In each case, the problem parameters are treated as random variables, and the aim is to learn the distribution of the solution. The results demonstrate that the generative PINN models can accurately capture both the mean and higher-order statistical properties of the solutions. Comparisons with reference solvers show very good qualitative and quantitative agreement, though challenges arise in representing sharply peaked distributions and boundary behaviours. Different architectures, including standard PINNs and PINN-PCE variants, are assessed in terms of convergence, error behaviour, and their effectiveness in representing uncertainty. Finally, Section 6 provides formal definitions and foundational elements used throughout the work.
@@
Physics-Informed Neural Networks (PINNs) are a class of neural network-based methods designed to approximate solutions of partial differential equations (PDEs), where the loss function incorporates the residual of the PDE [Karniadakis:pinn:orig:2019]. Similar approaches have been explored in [Lagaris_1998, Berg_2018, Raissi_2018, SSpiliopoulos:2018, makr_pryer2024deepUz],
To the best of our knowledge, the systematic framework presented here is novel. However, a number of related studies share certain aspects with our approach. A review of neural network approaches to uncertainty quantification can be found in [PSAROS2023111902]. A method which uses the PINN framework for random differential equations is presented as well. It is assumed the approximate solution is distributed with a normal distribution and the method provides ways to estimate its mean and variance. Inferencing stochastic solution and parameters of a Random PDE, by combining Polynomial Chaos Expansions, [PCE_karniadakis], and PCA we considered in [ZHANG2019108850, zhang2019learningmodalspacesolving], other related works include [ping2024uncertainty]. Physics Informed Generative Adversarial Networks were considered in [gao2023pi] and [yang2020physics]. See also [meng2022learning, MAL-056, unlu2020variational], for UQ related approaches. Neural Operators to address certain issues involving stochasticity in approximating PDEs in infinite dimensional setting include [lu2019deeponet, kovachki2023neural, Mishra_Conv_operNEURIPS2023, kovachki2024operator, li2020fourier, de2022cost, magnani2022approximate], approaches were proposed in [lu2021learning, Yang_2021, Liu_2023, figueres2025pinndomaindecomposition], their new role and importance in application to sciences was discussed in [koumouts_UQNN_review].
@@ 2.1 Problem and Assumptions
$\begin{split}&A_{\xi}u(x,t;\xi)=f(x,t;\xi)\quad\forall\left(x,t\right)\in D\times(0,T]\\ &B_{\xi}u(x,t;\xi)=g(x,t;\xi)\quad\forall\left(x,t\right)\in\Gamma\end{split}$ (6)
where $u(x,t;\xi)$ denotes the solution of the PDE, which we interpret as the mapping $U:\Xi\to\mathcal{H}$ with $U(\xi):=u(\cdot,\xi)$. Here, $A_{\xi}:\mathcal{H}\to\tilde{\mathcal{H}}$ is a $\xi-$parametrised differential operator, and $B_{\xi}:\mathcal{H}\to\overline{\mathcal{H}}$ is the associated boundary operator. The random parameter $\xi\in\Xi$ is distributed according to a Borel probability measure $\gamma$ on $\Xi$.
Analysing the behaviour of solutions is beyond the scope of this work. Our objective is to propose numerical algorithms for approximating the distribution associated with the solution of the problems at hand. Therefore, we assume that the solution to (6) exists and is sufficiently regular for our purposes. In particular, we assume throughout that
  • The operators $A_{\xi}$ and $B_{\xi}$ are well-defined for $\gamma-$almost every $\xi$, and the solution of (6) belongs in $L^{2}(\gamma;\,\mathcal{H})$.
@@
    $\begin{split}&\mathcal{A}\mathcal{Q}(\xi):=A_{\xi}q(\xi),\quad\mathcal{B}\mathcal{Q}(\xi):=B_{\xi}q(\xi)\qquad\text{ are }L^{2}-\text{integrable, i.e.,}\\[3.0pt] &\mathcal{A}\mathcal{Q}\in L^{2}(\gamma;\tilde{\mathcal{H}})\quad\text{ and }\quad\mathcal{B}\mathcal{Q}\in L^{2}(\gamma;\overline{\mathcal{H}})\,.\end{split}$ (7)
Where the space $L^{2}(\gamma;\mathcal{H})$ is defined in the sense of Bochner integration, see for example Section 7.2.2 of [Dashti_Stuart2017] for details. In particular this implies that
@@ 2.2 Probabilistic framework
We review some standard concepts from probability theory which will be useful in the sequel. We begin by assuming an abstract probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, a measurable space over $\mathcal{H}$, eg. $\left(\mathcal{H},\mathcal{B}\left(\mathcal{H}\right)\right)$, and a random element $\tilde{q}:\Omega\to\mathcal{H}$, ie. $\mathcal{F}/\mathcal{B}\left(\mathcal{H}\right)-$measurable function. The law of $\tilde{q}$ is the pushforward measure of $\mathbb{P}$ under $\tilde{q}$, defined as:
@@
Let $\xi:\Omega\to\Xi$ be a random element ( ie. $\mathcal{F}/\mathcal{B}\left(\Xi\right)-$measurable function) with law $\gamma$. We are interested in random elements $\tilde{q}$ expressible as compositions of a function $q\in L^{2}\left(\gamma;\,\mathcal{H}\right)$ and $\xi$, ie.
$\tilde{q}(\omega)=q\left(\xi(\omega)\right),\qquad\mathbb{P}-\text{as }\omega.$ (8)
@@
Hence setting $\nu_{q}:=\left[q\right]_{\!\#}\gamma$ we have
@@
where $\mathcal{P}(\mathcal{H})$ denotes the set of all Borel probability measures on $\mathcal{H}$, ie. measures defined on $\left(\mathcal{H},\mathcal{B}\left(\mathcal{H}\right)\right)$.
@@
    $\mathcal{P}_{p}\left(\mathcal{H}\right):=\left\{\mu\in\mathcal{P}\left(\mathcal{H}\right):\,\int_{\mathcal{H}}d_{\mathcal{H}}(h,\overline{h})^{p}\mu({\rm d}h)<\infty\quad\text{ for some (and hence any) }\overline{h}\in\tilde{\mathcal{H}}\right\}$
@@
Over those collections, one may consider alternative metrics for distances or divergences. In this work, we shall use the Wasserstein distance $W_{p}$, [AGS:2005, Villani_OT_book_2009] and the corresponding $\left(\mathcal{P}_{p}\left(\mathcal{H}\right),W_{p}\right)$ metric space. The $p-$Wasserstein distance is given by
$W_{p}(\mu,\nu):=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int_{\mathcal{H}\times\mathcal{H}}\,d_{\mathcal{H}}(x,y)^{p}\,\,\pi({\rm d}x,{\rm d}y)\right)^{1/p}\,,$ (9)
where $\Pi(\mu,\nu)$ is the set of all plans (joint distributions) with marginals $\mu$ and $\nu$, [AGS:2005]. Furthermore, we shall use the notation,
$d_{p,\gamma}(\left[X\right]_{\!\#}\gamma,\left[Y\right]_{\!\#}\gamma)^{p}:=\int_{\mathcal{H}\times\mathcal{H}}d_{\mathcal{H}}(x,y)^{p}\,\left[\left(X,Y\right)\right]_{\!\#}\gamma({\rm d}x,{\rm d}y)=\int_{\Xi}\left\|X(\xi)-Y(\xi)\right\|_{\mathcal{H}}^{p}\,\gamma({\rm d}\xi)\,.$ (10)
@@ 2.3 Transformed Law due to Parameterised Operators
We now introduce a notation to represent the law of expressions such as $A_{\xi}q(\xi)$ and $B_{\xi}q(\xi)$. In the most general setting, one might define stochastic processes $\tilde{A}:\Xi\times\mathcal{H}\to\tilde{\mathcal{H}}$ and $\tilde{B}:\Xi\times\mathcal{H}\to\overline{\mathcal{H}}$, with appropriate measurability conditions. However, in Section 2.1, we simplify the setting by assuming hypothesis stated in (7), ensuring measurability. We then define the operation $\odot$ as follows:
$A\odot\left(\left[q\right]_{\!\#}\gamma\right)(C):=\left[A_{\xi}q(\xi)\right]_{\!\#}\mathbb{P}(C)\,,$
@@
$A\odot\nu_{u}(C)=\nu_{f}(C)=\left[f(\cdot;\xi)\right]_{\!\#}\mathbb{P}(C),\qquad\forall C\in\mathcal{B}\left(\tilde{\mathcal{H}}\right)\,.$
It will be useful to summarise these definitions as follows. We define and study metric spaces of Borel probability measures, such as $\mathcal{P}_{2}^{\gamma}(\mathcal{H}),\,\mathcal{P}_{2}^{\gamma}(\tilde{\mathcal{H}})$ and $\mathcal{P}_{2}^{\gamma}(\overline{\mathcal{H}})$, and decompose the law of the solution ${u}$ into a measurable function $U(\xi)$ and a probability measure $\gamma$, according to the diagram:

### v2

# Neural Measures for Learning Distributions of Random PDEs

@@ Abstract
keywords:
Random PDEs, Neural networks for PDEs, Generative models, Neural measure, Physics-Informed Neural Networks, Universal Approximation Theorem, Uncertainty Quantification.
{MSCcodes}
65C20, 60H35, 35R60
@@ 1.1 Problem Formulation and Approach
(1) $\begin{split}&\mathscr{A}_{\xi}u(\cdot,\xi)=\mathscr{F}_{\xi},\end{split}$
where $\mathscr{A}_{\xi}$ is a $\xi$-parametrised differential operator (incorporating possible initial and boundary conditions), and $\mathscr{F}_{\xi}$ is the source. For each fixed $\xi$, $u(\cdot;\xi)$ denotes the solution of the differential equation, which we interpret as a mapping $U:\Xi\to\mathcal{H}$ with $U(\xi)\coloneqq u(\cdot,\xi)$. The space $\mathcal{H}$ where $u$ takes values for each $\xi$ is an (infinite dimensional) Hilbert or Sobolev space. We assume that the random parameter $\xi\in\Xi$ is distributed according to a known Borel probability measure $\gamma$ on $\Xi$. Our aim is to approximate the random variable $U$ and its law $\nu_{u}\in\mathcal{P}(\mathcal{H})\,;$ here $\mathcal{P}(\mathscr{X})$ denotes the set of Borel probability measures on $\mathscr{X}$.
@@
One of the key successes of neural network methods is their ability to effectively approximate probability distributions. In fact, several important problems of interest rely on generating a probability distribution when the available data represent specific instances of the unknown distribution. Typically, if the data is represented by an empirical measure $\nu_{\,\text{data}}\in\mathcal{P}(\mathbb{R}^{k})$, we would like to generate an approximation of $\nu_{\,\text{true}}\in\mathcal{P}(\mathbb{R}^{k})$ by optimising
(2) $\min_{\mu\in V_{\Theta}}d_{\mathscr{M}}\left(\mu,\nu_{\,\text{data}}\right),$
@@
(3) $\mu_{\theta^{*}}\cong\nu_{\,\text{true}},\qquad\text{where }\mu_{\theta^{*}}\in\operatorname*{arg\,min}_{\mu\in V_{\Theta}}d_{\mathscr{M}}\left(\mu,\nu_{\,\text{data}}\right).$
Representative works utilising the aforementioned generative model framework for a variety of applications include [1, 17, 4, 59, 9, 19], and the references in [55]. For a recent modelling perspective that interprets various generative models as mean field games, see [55]. Recent advances connect random PDEs with foundational models [56].
The above framework relies on various alternative choices, each one yielding fundamentally distinct algorithms at the end. We would like to explore the potential extension of this approach to approximate solutions of Eq. 1. To accomplish this we must first systematically construct discrete spaces $V_{\Theta}$ such that $V_{\Theta}\subset\mathcal{P}(\mathcal{H}).$ We adopt the following alternative approaches:
@@
Furthermore, we show that the neural measure spaces are expressive enough to approximate appropriate subspaces of $\mathcal{P}(\mathcal{H})$.
@@
(4) $\mathscr{A}\odot\mu,\qquad\mu\in V_{\Theta}\subset\mathcal{P}(\mathcal{H}),$
@@
(5) $\min_{\mu\in V_{\Theta}}d_{\tilde{\mathscr{M}}}\left(\mathscr{A}\odot\mu,\nu_{\,\mathscr{F}}\right),$
[…]
@@
This work is structured to systematically develop and analyse a generative modelling approach for Uncertainty Quantification using Physics-Informed Neural Networks (PINNs). In Section 2, we introduce the underlying problem, and the corresponding probabilistic framework essential to our study. Next, we introduce new notational conventions and the variational formulation used. The variational principle forms the backbone of our approach, guiding the optimisation process for learning the solution distribution of the random differential equation. Section 3 presents the model construction in detail. We define a fundamental representation of Neural Measures tailored to the problem’s probabilistic nature and explore three distinct architectural frameworks, namely fully discrete randomised neural networks, Polynomial Chaos Expansion Neural Networks (PCE-NN), Galerkin-based neural networks. In Section 4, we shift focus to the variational distances used during training, with particular emphasis on the Wasserstein distance and its interpretation in terms of Dirac measures, which naturally relate it to $L^{p}$-type loss functionals. We also discuss alternative metrics and divergences that can be integrated into the framework. Section 5 presents numerical experiments that validate the proposed framework across three representative problems: a bistable ODE, a linear diffusion PDE, and a reaction-diffusion PDE. In each case, the problem parameters are treated as random variables, and the aim is to learn the distribution of the solution. The results demonstrate that the generative PINN models can accurately capture both the mean and higher-order statistical properties of the solutions. Comparisons with reference solvers show very good qualitative and quantitative agreement, though challenges arise in representing sharply peaked distributions and boundary behaviours. Different architectures, including standard PINNs and PINN-PCE variants, are assessed in terms of convergence, error behaviour, and their effectiveness in representing uncertainty. Appendix A presents universal approximation theorems showing that the neural measure spaces, suggested herein, are expressive enough. Therein, we also provide the formal definitions and foundational elements used throughout the work.
@@
Physics-Informed Neural Networks (PINNs) are a class of neural network-based methods designed to approximate solutions of partial differential equations (PDEs), where the loss function incorporates the residual of the PDE [41]. Similar approaches have been explored in [24, 5, 40, 45, 31].
To the best of our knowledge, the systematic framework presented here is novel. However, a number of related studies share certain aspects with our approach. A review of neural network approaches to uncertainty quantification can be found in [39]. A method which uses the PINN framework for random differential equations is presented as well. It is assumed the approximate solution is distributed with a normal distribution and the method provides ways to estimate its mean and variance. Inferencing stochastic solution and parameters of a Random PDE, by combining Polynomial Chaos Expansions, [50], and PCA we considered in [58, 57], other related works include [37]. Physics Informed Generative Adversarial Networks were considered in [16] and [53]. See also [32, 21, 48], for UQ related approaches. Neural Operators to address certain issues involving stochasticity in approximating PDEs in infinite dimensional setting include [28, 22, 42, 23, 26, 12, 30], approaches were proposed in [29, 52, 27, 14], their new role and importance in application to sciences was discussed in [15].
@@ 2.1 Problem and Assumptions
(6) $\begin{split}&A_{\xi}u(x,t;\xi)=f(x,t;\xi)\quad\forall\left(x,t\right)\in D\times(0,T]\\ &B_{\xi}u(x,t;\xi)=g(x,t;\xi)\quad\forall\left(x,t\right)\in\Gamma\end{split}$
where $u(x,t;\xi)$ denotes the solution of the PDE, which we interpret as the mapping $U:\Xi\to\mathcal{H}$ with $U(\xi)\coloneqq u(\cdot,\xi)$. Here, $A_{\xi}:\mathcal{H}\to\tilde{\mathcal{H}}$ is a $\xi$-parametrised differential operator, and $B_{\xi}:\mathcal{H}\to\overline{\mathcal{H}}$ is the associated boundary operator. The random parameter $\xi\in\Xi$ is distributed according to a Borel probability measure $\gamma$ on $\Xi$.
Analysing the behaviour of solutions is beyond the scope of this work. Our objective is to propose numerical algorithms for approximating the distribution associated with the solution of the problems at hand. Therefore, we assume that the solution to Eq. 6 exists and is sufficiently regular for our purposes. In particular, we assume throughout that
  • The operators $A_{\xi}$ and $B_{\xi}$ are well-defined for $\gamma$-almost every $\xi$, and the solution of Eq. 6 belongs in $L^{2}(\gamma;\,\mathcal{H})$.
@@
    (7) $\begin{split}&\mathcal{A}Q(\xi)\coloneqq A_{\xi}q(\xi),\quad\mathcal{A}Q(\xi)\coloneqq B_{\xi}q(\xi)\qquad\text{ are }L^{2}-\text{integrable, i.e., }\\[3.0pt] &\mathcal{A}Q\in L^{2}(\gamma;\tilde{\mathcal{H}})\quad\text{ and }\quad\mathcal{B}Q\in L^{2}(\gamma;\overline{\mathcal{H}}).\end{split}$
Where the space $L^{2}(\gamma;\mathcal{H})$ is defined in the sense of Bochner integration, see for example Section 7.2.2 of [11] for details. In particular this implies that
@@ 2.2 Probabilistic framework
We review some standard concepts from probability theory which will be useful in the sequel. We begin by assuming an abstract probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$, a measurable space over $\mathcal{H}$, eg. $\left(\mathcal{H},\mathcal{B}\left(\mathcal{H}\right)\right)$, and a random element $\tilde{q}:\Omega\to\mathcal{H}$, i.e., $\mathcal{F}/\mathcal{B}\left(\mathcal{H}\right)$-measurable function. The law of $\tilde{q}$ is the pushforward measure of $\mathbb{P}$ under $\tilde{q}$, defined as:
@@
Let $\xi:\Omega\to\Xi$ be a random element ( i.e., $\mathcal{F}/\mathcal{B}\left(\Xi\right)$-measurable function) with law $\gamma$. We are interested in random elements $\tilde{q}$ expressible as compositions of a function $q\in L^{2}\left(\gamma;\,\mathcal{H}\right)$ and $\xi$, i.e.,
(8) $\tilde{q}(\omega)=q\left(\xi(\omega)\right),\qquad\mathbb{P}-\text{as }\omega.$
@@
Hence setting $\nu_{q}\coloneqq\left[q\right]_{\!\#}\gamma$ we have
@@
where $\mathcal{P}(\mathcal{H})$ denotes the set of all Borel probability measures on $\mathcal{H}$, i.e., measures defined on $\left(\mathcal{H},\mathcal{B}\left(\mathcal{H}\right)\right)$.
@@
    $\mathcal{P}_{p}\left(\mathcal{H}\right)\coloneqq\left\{\begin{split}\mu\in\mathcal{P}\left(\mathcal{H}\right):\,\int_{\mathcal{H}}d_{\mathcal{H}}(h,\overline{h})&{}^{p}\mu({\rm d}h)<\infty\\ &\text{for some (and hence any) }\overline{h}\in\tilde{\mathcal{H}}\end{split}\right\}$,
@@
Over those collections, one may consider alternative metrics for distances or divergences. In this work, we shall use the Wasserstein distance $W_{p}$, [2, 49] and the corresponding $\left(\mathcal{P}_{p}\left(\mathcal{H}\right),W_{p}\right)$ metric space. The $p$-Wasserstein distance is given by
(9) $W_{p}(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\mu,\nu)}\left(\int_{\mathcal{H}\times\mathcal{H}}\,d_{\mathcal{H}}(x,y)^{p}\,\,\pi({\rm d}x,{\rm d}y)\right)^{1/p},$
where $\Pi(\mu,\nu)$ is the set of all plans (joint distributions) with marginals $\mu$ and $\nu$, [2]. Furthermore, we shall use the notation,
(10) $\begin{split}d_{p,\gamma}(\left[X\right]_{\!\#}\gamma,\left[Y\right]_{\!\#}\gamma)^{p}:&=\int_{\mathcal{H}\times\mathcal{H}}d_{\mathcal{H}}(x,y)^{p}\,\left[\left(X,Y\right)\right]_{\!\#}\gamma({\rm d}x,{\rm d}y)\\ &=\int_{\Xi}\left\|X(\xi)-Y(\xi)\right\|_{\mathcal{H}}^{p}\,\gamma({\rm d}\xi).\end{split}$
@@ 2.3 Transformed Law due to Parameterised Operators
We now introduce a notation to represent the law of expressions such as $A_{\xi}q(\xi)$ and $B_{\xi}q(\xi)$. In the most general setting, one might define stochastic processes $\tilde{A}:\Xi\times\mathcal{H}\to\tilde{\mathcal{H}}$ and $\tilde{B}:\Xi\times\mathcal{H}\to\overline{\mathcal{H}}$, with appropriate measurability conditions. However, in Section 2.1, we simplify the setting by assuming hypothesis stated in Eq. 7, ensuring measurability. We then define the operation $\odot$ as follows:
$A\odot\left(\left[q\right]_{\!\#}\gamma\right)(C)\coloneqq\left[A_{\xi}q(\xi)\right]_{\!\#}\mathbb{P}(C),$
@@
[…]
$\mathbb{P}\quad\xrightarrow{\xi}\quad\gamma\quad\begin{cases}\xrightarrow{q}\quad\left[q\right]_{\!\#}\gamma\\ \xrightarrow{A\odot\,}\quad A\odot\,\left[q\right]_{\!\#}\gamma\\ \xrightarrow{B\odot\,}\quad B\odot\,\left[q\right]_{\!\#}\gamma,\end{cases}$